An appliance dealer sells three different models of upright freezers having 13.3, 16.1, and 19.3 cubic feet of storage space. Consider the random variable x = the amount of storage space purchased by the next customer to buy a freezer. Suppose that x has the following probability distribution.

x p(x)
13.3 0.2
16.1 0.5
19.3 0.3

(a)

Calculate the mean and standard deviation of x (in cubic feet). (Hint: See Example 6.15. Round your standard deviation to four decimal places.)

μx= cubic ft

σx= cubic ft
Answer #1

Solution :

x P(x) x * P(x) x2 * P(x)
13.3 0.2 2.66 35.378
16.1 0.5 8.05 129.605
19.3 0.3 5.79 111.747
Sum 1 16.5 276.73

μx = X * P(X) = 16.5 cubic ft

σx = X 2 * P(X) - 2 = 276.73 - 16.52 = 2.1166 cubic ft